OpenStudy (anonymous):

Greg has a piece of rope that is 18 ft long. He wants to cut it into two pieces so that the longer piece is 2 ft less than 3 times the length of the shorter piece. How long will each piece of rope be? Shorter piece: ft Longer piece: ft

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@Data_LG2 @Destinymasha

OpenStudy (anonymous):

so total length of rope is 18 ft. cut in half: let x be the shorter piece, so long piece would be : 2- 3x so you'll have long piece + short piece =total length of rope 2-3x + x =18 solve for x, tell me what you got and i'll tell you the rest

OpenStudy (anonymous):

x=-8

OpenStudy (anonymous):

oh crap sorry i made a mistake. longer piece should be 3x-2 not 2-3x. sorry so that will be 3x-2+x=18, x= 5 so since x is the length of the short piece then the short piece would be 5 ft long. now the longer piece is 3x-2, substitute 5 and you'll get?

OpenStudy (anonymous):

not for sure

OpenStudy (anonymous):

3(5)-2?

OpenStudy (anonymous):

oh ok the x confused me sorry

OpenStudy (anonymous):

13?

OpenStudy (anonymous):

x- is the shorter piece of the rope.. i just named the length of the shorter piece as "x"

OpenStudy (anonymous):

long piece = 2 ft less than 3 times the length of the shorter piece if i let the short piece equal to x, then the longer piece is will be 2 ft less than 3 times the length of the "x". the mathematical interpretation will be longer piece is 3x - 2 is it clear?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

will you help me with another one

OpenStudy (anonymous):

Admission to the museum is $12 for children and $15 for adults. On a certain day, 2450 people go to the museum and $33,456 is collected. How many children and how many adults went to the museum? children__ adults__

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

let x - children let y-adult number of people attended x+y = 2450 cost 12x + 15y = 33,456 so you have TWO equations and you have to solve for TWO variables. do you know how to solve systems of linear equations?

OpenStudy (anonymous):

no